# KineSoft: Learning Proprioceptive Manipulation Policies with Soft Robot Hands

### v1

@@ V-A Baselines
Shape-tracking Baselines. For shape tracking and task performance evaluation we provide the results against the following:
@@
  • Strain Policy. Strain policy, based on prior works that directly use sensor readings without intermediate representations for learning manipulation policies [30], uses raw sensor measurements instead of reconstructed shapes. For shape tracking evaluation, we modified the low-level controller from Section IV-D to track reference sensor readings directly through proportional tendon actuation. For task performance evaluation, we trained a diffusion policy using the same 50 demonstrations we use for KineSoft, but with raw strain signals and wrist-mounted camera observations as input states.
TABLE I: Shape Tracking Error Comparison
Method Representation Error (mm)
Strain-tracking [6, 30] Strain $6.20\pm 2.39$
KineSoft Mesh $\mathbf{3.29\pm 0.91}$
Figure 6: Trajectory tracking performance evaluation. We provide each of the controllers with kinesthetically deformed shape trajectories.
Figure 7: Tasks and KineSoft Rollout.
V-B Shape Estimation Fidelity
We evaluated the proposed shape estimation model against two baseline methods from the literature: constant curvature model [9, 45], a common analytical approach for soft robot shape representation, and DeepSoRo [38], a learning-based point cloud reconstruction method. Additionally, we demonstrate that our soft body mechanics-based mapping between observed resistance values and simulated strains, as outlined in Section IV-B, substantially improves upon a naively calibrated linear mapping while using the same domain alignment approach proposed in Section IV-C.
We performed evaluations using observed point clouds generated by the physical setup shown in Figure 2, using the unidirectional Chamfer distance metric as outlined in Equation 7. The quantitative results are outlined in Table II and visualized in Figure 5. Our model-based approach achieved a shape estimation error of $1.92$ mm, representing a 41.3% improvement over the best baseline method (DeepSoRo) and a 60.9% improvement over the linear variant of our approach.
TABLE II: Shape Estimation Fidelity
Method Input Signals Representation Shape Error [mm]
PneuFlex Sensor Strain curvature $3.70\pm 1.36$
DeepSoRo Strain point cloud $3.27\pm 1.05$
KineSoft Strain (naive) mesh $4.91\pm 2.85$
KineSoft Strain (model) mesh $\mathbf{1.92\pm 0.39}$
V-C Shape-conditioned Controller Performance
We evaluated our shape-conditioned controller against a strain-tracking baseline that directly uses sensor readings for control, as commonly implemented in prior soft robot manipulation works [6, 30]. For evaluation, we collected reference trajectories through kinesthetic teaching, where a demonstrator physically deformed the fingers into desired configurations. During execution, the controllers had to track these trajectories using tendon actuation.
This evaluation specifically highlights a fundamental challenge in soft robot imitation: the sensor signals generated during kinesthetic demonstration (when fingers are manually deformed) differ significantly from those produced during autonomous execution (when tendons are actuated). The strain-tracking baseline, which attempts to directly match these sensor readings, struggles with this demonstration-execution gap, achieving $6.20$ mm tracking error. In contrast, our shape-conditioned controller bridges this gap by tracking the geometric shape itself rather than the underlying sensor signals, achieving $3.29$ mm error, which is a 47% improvement.
Figure 4 visualizes this performance difference. While both controllers receive reference trajectories collected through kinesthetic teaching, the shape-conditioned controller successfully reproduces the demonstrated shapes using tendon actuation, even though the sensor signals during execution differ from those during the demonstration. The strain-tracking baseline exhibits larger errors because it attempts to match sensor readings that are different between demonstration and execution modes.
The improved shape tracking performance of KineSoft’s shape-conditioned controller stems from its ability to abstract away the differences in how deformations are achieved (manual guidance vs. tendon actuation) by focusing on the resulting geometric configurations. This shape-based abstraction provides a consistent representation across demonstration and execution, enabling functionally successful reproduction of demonstrated skills.
V-D In-hand Manipulation Task Performance
In-hand Manipulation Tasks. We evaluated KineSoft on in-hand manipulation tasks requiring precise finger control and in-hand dexterity:
  • Cap Unscrewing. MOE makes contact with a bottle cap that is screwed on. Then it must coordinate the three fingers to rotate the cap sufficiently to unscrew it then raise it up. The task is considered successful if the initially screwed on container is detached from the bottle.
  • Container Unlidding. MOE initially grasps an empty container that has a closed hinged snap top lid. It must then lift up the container then move the fingers to the right positions and flip the lid open. To apply sufficient force on the lid, two of the MOE fingers must hold the container’s body while one of the fingers pries open the lid. The task is considered successful if the container lid is open in MOE’s grasp.
For each task, we collected 50 successful demonstrations through kinesthetic teaching, where an expert demonstrator physically guided the fingers through the required motions. This kinesthetic teaching approach highlights a key advantage of our framework: while humans can intuitively demonstrate complex manipulations by directly deforming the soft fingers, these demonstrated configurations must ultimately be achieved through tendon actuation during autonomous execution. This creates a gap, as the sensor signals during human demonstration (direct finger deformation) differ substantially from those during autonomous execution (tendon actuation). KineSoft bridges this gap by focusing on reproducing the demonstrated shapes rather than matching the sensor signals directly.
As shown in Table III, KineSoft achieved a 100% success rate (5/5 trials) on the unscrewing task compared to 0/5 for the strain-matching baseline. The baseline policy failed primarily because it attempted to reproduce sensor signals from demonstration that were either impossible to achieve through tendon actuation alone or not grounded to the functionally correct robot states to perform the task. In contrast, KineSoft’s shape-based approach successfully translated the demonstrated trajectories into executable tendon commands that produced the desired finger configurations.
These results demonstrate that by focusing on tracking demonstrated shapes rather than raw sensor signals, KineSoft effectively bridges the gap between human demonstration and autonomous execution. The shape estimation model provides a consistent representation across both modes, while the shape-conditioned controller reliably reproduces the demonstrated behaviors despite the gap in how deformations are achieved.
TABLE III: Task Performance Across Different Policies
Task Policy State Representation Success Rate
Cap Unscrewing Strain-Matching Strain 0/5
KineSoft Mesh 5/5
Container Unlidding Strain-Matching Strain 0/5
KineSoft Mesh 4/5
VI Limitations
While KineSoft demonstrates effective shape estimation and trajectory tracking for soft robotic manipulation, several limitations remain. First, the current system lacks explicit force feedback mechanisms, relying instead on kinematic trajectories and deformation states to encode manipulation skills. This limits precise force control during object interactions.
Additionally, although kinesthetic teaching provides an intuitive approach to demonstrating some behaviors, it requires the expert demonstrator to recognize the MOE finger’s workspace limitations to avoid demonstrating shapes that the MOE cannot reach. In our future work, we aim to leverage KineSoft’s real-time shape estimation to communicate robot affordances to users via a visual user interface during demonstrations.
Although the sensors used in this work are readily available and low-cost, they introduce fabrication complexity when integrating into the prior MOE design, which was relatively simple to fabricate. We also observed that the sensors may not respond dynamically in fast manipulation tasks due to commonly observed soft body phenomena such as creep and hysteresis, where the sensor signals require some time to reach equilibrium conditions when stretched quickly.
The framework also requires careful calibration of the strain sensors and domain alignment parameters to achieve reliable shape estimation. To this end, the proposed domain alignment method provides an algorithmic approach to self-calibrate and align the real-world sensor readings. However, this process still requires a controlled test environment, as shown in Figure 2, which may limit reproducibility and accessibility.
KineSoft’s shape-conditioned low-level controller weighs the residuals from vertices of the reference meshes equally. Implicitly, the controller prioritizes the vertices closer to the tip higher because the displacement tends to be greater further away from the base. For many tasks, this is an appropriate weighing as the finger tips play a vital role in performing dexterous in-hand manipulation tasks [7]. However, different tasks may require different regions of the fingers to be prioritized based on contact probability, for instance.
VII Conclusion and Lessons
This paper presents KineSoft, a framework for learning dexterous manipulation skills with soft robot hands that embraces, rather than fights against, their inherent compliance. The key insight of our work is recognizing that while this compliance enables intuitive kinesthetic teaching, it creates a fundamental gap between demonstration and execution, where the deformations and sensor signals during human demonstration differ substantially from those during autonomous execution through tendon actuation.
KineSoft addresses this challenge through a hierarchical approach. A shape estimation model provides consistent geometric representations across demonstration and execution modes while a domain alignment method enables robust transfer of simulation-trained models to real hardware and a shape-conditioned controller reliably tracks the policy’s generated deformation trajectories despite the different underlying actuation mechanisms. Then, a high-level imitation policy learns to generate target vertex deformations from demonstrations, capturing the intended manipulation strategy in geometric grounding. Our experimental results demonstrate that this shape-based hierarchical approach enables more effective skill transfer than methods that attempt to directly match sensor signals or joint configurations. This work suggests that successful imitation learning for dexterous soft robots requires careful consideration of how demonstration and execution modes differ, and appropriate intermediate representations to bridge this gap.
Appendix A Sensor Signals
@@
Appendix B Other Task
Figure 9: Strain tracking performance on a simple task
Strain tracking can succeed in some tasks. Particularly with a weighted bottle opening task, where one finger could flick the lid open, the strain matching policy seemed to perform consistently.

### v2

@@ Appendix D Baselines
Shape-tracking Baselines. For shape tracking and task performance evaluation we provide the results against the following: Strain Policy: Strain policy, based on prior works that directly use sensor readings without intermediate representations for learning manipulation policies [51], uses raw sensor measurements instead of reconstructed shapes. For shape tracking evaluation, we modified the low-level controller from Section 4.4 to track reference sensor readings directly through proportional tendon actuation. For task performance evaluation, we trained a diffusion policy using the same 50 demonstrations we use for KineSoft, but with raw strain signals and wrist-mounted camera observations as input states.
Appendix E Data Generation and Shape Estimation Model Training
To train the model, we generate a large dataset of deformed meshes using SOFA (Simulation Open Framework Architecture) [60]. We simulate a tetrahedral finite-element mesh of the MOE finger with a Neo-Hookean hyperelastic material model parameterized by elastic material properties that are randomized at runtime.
We model the tendon actuation with massless, inextensible cables running through a series of fixed points within the finger body. We discretize each tendon path to segments defined by 3D attachment points embedded in the tetrahedral mesh. The cable constraint applies forces to these points to maintain constant length while allowing sliding, effectively simulating the mechanical behavior of Bowden cable transmission. The soft body scene is solved with an implicit Euler time integration scheme and uses a conjugate gradient solver for the system matrices. We generate training data by randomly sampling tendon actuation commands within the feasible range and recording the resulting deformed vertex positions and embedded sensor strains. To simulate rich deformation behaviors including contact-like effects, we apply random external forces to the finger surface. These forces are randomly applied over time with sufficiently large radii to ensure smooth deformations that mimic natural contact interactions, without requiring explicit and difficult-to-model contacts in the scene.
We train the model using a mean squared error (MSE) loss on vertex displacements:
where $\Delta\mathbf{v}_{j,i}^{*}$ represents the ground-truth displacement for vertex $i$ of mesh finger $j$. This choice of loss function provides strong supervision by enforcing explicit vertex-wise correspondence between predicted and ground-truth meshes. Because we leverage simulated data to train the model, we can exploit the vertex-level correspondences in the meshes. We contrast this to prior works that relied on chamfer distance loss over real-world partial observations [50], MSE loss ensures that each vertex learns to track its specific local deformation patterns, enabling precise reconstruction of the full finger shape.
Appendix F Experiment Setup
Figure 7: Simulation and real-world setup. A: Simulated robot workspace and sample of simulated strain signals. B: Real-world robot setup with the projected patterns to improve ground-truth shape observations for evaluation and calibration.
Appendix G Failiure Cases
We conducted stress testing through 1,000 continuous servo cycles over 14 hours, observing maximum signal drift of 8.3% in a sensor (Fig. LABEL:fig:rebuttalA) that resulted in $0.062\pm 0.0056$ mm mean vertex error. After rest, the signals returned to baseline levels and our original manipulation experiments were conducted over several weeks without recalibration. We appreciate the reviewers highlighting this important concern and will update the manuscript with these discussions on long-term sensor reliability.
KineSoft failures occurred mainly when fingers lost or made unintended contact during manipulation transitions—a limitation of kinematic imitation without tactile sensing, which we will highlight in the limitations section. Baseline strain policies failed significantly more, as demonstration strain patterns often cannot be directly reproduced through tendon actuation. The results support KineSoft’s approach of using shape as an intermediate representation between demonstration and execution modes.
Figure 8: Left: Sensor signal degradation after 1000 cycles. Right: Failiure cases for manipulation tasks.
Appendix H Sensor Signals